OpenStudy (anonymous):

You have a pool that measures 15 yards by 20 yards. You want to build a deck around all four sides of the pool, and you want to increase the total area to 600 square yards. What is the width of the deck? aproximately 3.6 yards Approximately 4.5 yards Approximately 5.1 yards Approximately 6.7 yards

OpenStudy (anonymous):

please help :(

OpenStudy (taramgrant0543664):

Well to start you can find the area of the pool A=(15)(20)=? The total area includes both the pool and the deck so that means you can take the total area (600) and subtract the area of the pool to get the area that the deck needs to take So this is how I would set it up hopefully my not so lovely drawing gets included: So you will have 4 squares which each can be represented by A=x^2 We also have the top and bottom which each can be represented by A=20x And we also have the two sides which each can be represented by A=15x Keeping in mind that the area of the deck should be 300 since the area of the pool is 300 To put this all together now: 300=4x^2+2(20x)+2(15x) Simplifying we get: 300=4x^2+40x+30x 300=4x^2+70x Factoring out 2 we get: 150=2x^2+35x Soon you can put your options for the width in to determine which one it is

OpenStudy (anonymous):

where would I plug those options in at?

OpenStudy (taramgrant0543664):

150=2x^2+35x

OpenStudy (anonymous):

so into the x's?

OpenStudy (taramgrant0543664):

Yes the options that you have above would be the x's